You have two saltwater solutions. One consists of 20% salt, and the other consists of 15% salt. If you wanted to make 5 gallons of a 19% solution, what volume of each solution should you add?

Answer by checkley79(3341)

.20X+.15(5-X)=.19*5
.20X+.75-.15X=.85
.10X=.95-.75
.10X=.20
X=.20/.10
X=2 GAL OF THE 20% MIX.
.20*2+.15(5-2)=.19*5
.40+.15*3=.85
.40+.45=.85
.85=.85